Complete the the given ordered pairs for the queations and graph the equations: What are the correct steps to solve. Help
X-Y=3
(0__,),(___,0)
(5,__),(2,___)
6x-5y=30
(o,__),(___,0)
(3,__),(___,-2)

Answer by Earlsdon(6294)

Complete the ordered pairs for:
x-y+=+3
(0, __), (__, 0)
Substitute x = 0, then solve for y: y = -3
Substitute y = 0, then solve for x: x = 3, so...
(0, -3)
(3, 0)
----------
(5,__), (2,__)
Substitute x = 5, then solve for y: 5-y = 3, so y = 2
Substitute x = 2, then solve for y: 2-y = 3, so y = -1
(5, 2)
(2, -1)
-----------
6x-5y = 30
(0,__), (__,0)
Substitute x = 0, then solve for y: 6(0)-5y = 30, so y = -6
Substitute y = 0, then solve for y: 6x-5(0) = 30, so x = 5
(0, -6)
(5, 0)
----------
(3,__), (__, -2)
Substitute x = 3, then solve for y: 6(3)-5y = 30, 18-5y = 30, -5y = 12, y = -12/5 = -2.4
Substitute y = -2, then solve for x: 6x-5(-2) = 30, 6x+10 = 30, 6x = 20, x = 20/6 = 3.33
(3, -2.4)
(3.33, -2)